OpenStudy (anonymous):

The volume of a cylinder is 980in3 .The height of the cylinder is 20 in. What is the radius of the cylinder? Express your answer to the nearest hundredth.

jimthompson5910 (jim_thompson5910):

hint: use the formula \[\Large V = \pi*r^2*h\] where V = volume of cylinder \(\Large \pi\) = 3.14 (approximately) r = radius h = height

OpenStudy (anonymous):